# An Industry Interview Study of Software Signing for Supply Chain Security

### v1

@@ 5.1.2 Build
Next, we extract from our subjects’ responses how signing is used in the build process of software products. For most of the subjects, the build and build signing (for subjects who do so) are automated. Our model expects that the source code to be built should be verified (VSC), and then when the build is successful we expect another round of provenance certification (PB).
Verification of source code signatures before build (VB): Signatures from source codes are rarely verified before the commencement of the build process. Two subjects mentioned some sort of rigorous verification at this point. Alternate means of security at this stage are similar to the security techniques reported for the build process.
Signing of build output (PB): 13 subjects reported the use of signing at this stage of the software delivery process. 10 subjects reported signing is used alongside attestations, SBOMS, and other forms of metadata.
@@ 5.1.3 Package/Artifact
Verification of build signature before deployment of final artifact (VB): Similar to the other verification steps, signatures from the build phase are rarely verified before the deployment. Only two subjects mentioned verification at this point. For example, S7 states, “We sign our code and then we also verify that it was built in our specific build system”.
Other subjects mentioned that the security of the entire build process comprises techniques like an automated build pipeline with Workflow definitions, use of build policies, and Reproducible builds. These are complemented by code scanning tools as in the case of the source code stage.
Signing of Final Software product (PA): All 18 subjects said that signing is done at the deployment stage to establish provenance for the organization on each of their products.
@@ 5.1.4 Third-Party Dependencies
For third-party dependencies in the Software Supply Chain, we evaluated how software signing is used as a factor influencing dependency selection, and how software signing is used to verify the authenticity of dependencies in practice.
@@
Presence of a Signature as a factor for Dependency Selection: For most of our subjects, signatures on a third-party dependency do not affect their choice in selecting such a dependency. Only 3 of our subjects like to see this, treating signatures as a signal that the owner is security conscious. Other security factors that influence the selection of third-party dependencies as mentioned are: the use of the OpenSSF’s scorecard tool, the dependency’s known vulnerabilities, level of activity, licensing, and age.
@@
Verification of Signatures from external dependencies (VD): Signatures on third-party dependencies are rarely verified in practice. Only three subjects reported some form of verifying signatures from third-party dependencies. As S11 said, “If it is open source, then probably no…But if it is commercial off the shelf, we want to make sure those are signed.” However, most of our subjects reported several methods for assessing the authenticity of a third-party dependency. These methods include code and vulnerability scanning (e.g., Coverity), source rebuilding, internal mirroring of dependencies, and reliance on ecosystem internal checks (e.g., Go package manager’s checks). Other subjects perform internal reviews of dependencies, while some have a centralized dependency selection and management process, where dependencies are hosted and maintained by the organization. A few subjects noted that their organization discourages or forbids third-party dependencies.
@@
Signing after verification and certification of dependencies(PD): A few subjects reported that third-party dependencies are thoroughly reviewed using one of the aforementioned techniques before being signed off for use by a team member. Six of our subjects indicated that this process is followed in their software production workflow. For example, S17 described their team’s process like this: “We have a signed report that’s signed by us, and we trust us…that says, ‘Hey, I scanned this third-party dependency, did all these things. I don’t have a signature for that, but I ran all these extra checks and then signed those checks. So now we have a record”’. In contrast to that attestation approach, S18’s team ingests third-party dependencies and signs the internal versions directly.
@@ 5.2 RQ2: Challenges affecting the implementation of software signing in practice
Finding 2. Subjects identified several common themes as challenges faced in practice as obstacles to effectively implementing software signing. We categorize these broadly; 1. Technical Challenges: These include – key Management and Infrastructure issues, compatibility issues with existing systems and different architectures, and lack of signature verification 2. Organizational and human Challenges: These include – expertise in setting up Signing implementation, operationalization of the signing procedure, developer attitude, and Resources to set up the signing process
Our next research objective was to understand the various challenges practitioners face while implementing software signing in practice. We grouped subjects’ responses under three main themes for challenges faced by practitioners in practice; Technical, Organizational, and Human factors. Table IV gives a summary. Due to space limitations, we focus on only a subset of the individual challenges.
TABLE IV: Challenges to software signing Implementation in practice. We have categorized related challenges highlighted by subjects. Two uncommon Technical challenges are described in Appendix E.
Observed Challenges #Subjects(#Orgs) Proposed Solutions
@@
Key management 10 (9) Use of Keyless Signing (e.g., Sigstore)
Compatibility issues 6 (6)
Lack of Verification of Signatures 6 (5) Signed Metadata, Component Data Management
Ease of use/Usability 4 (4) Usable signing tools (e.g., Sigstore), Documentation
No unifying Standard 2 (2)
Lack of Verification of Signatures 6 (5) Signed Metadata, Component Data Management
@@
Operationalization of the Signing process 4 (4) Automating Signing
Resource to Set up Signing 3 (3)
Creating Effective Signing Policy 2 (2) Regular Process Feedback mechanisms
No Management incentive to Sign 2 (2)
Bureaucracy 1 (1)
@@
Expertise in setting setup and use of signing 5 (4)
Developer Attitude to Signing 3 (3) Automating Signing
Lack of Demand from Customers 1 (1)
@@ 5.2.1 Technical Challenges
Some subjects discussed recent signing technologies that implement keyless signing (e.g., Sigstore) as a possible solution. S18 said: “Sigstore comes in…keyless signing…you don’t have to worry about long-lasting static SSH keys, or keys being compromised, or rotating keys…you don’t have these big servers to store all these signatures”. S17 also noted, “[Sigstore’s] other strength…I can associate my identity with an OIDC identity as opposed to necessarily needing to generate a key and keep track of that key and yada, yada. So that’s super useful because I could say, ‘Oh, this is signed with [a] GitHub identity.’ So unless my GitHub identity has been compromised, that’s much better.”.
@@
Compatibility issues: Compatibility issues arise during the integration of software signing systems and tools with other tools within the software engineering process, including other security tools and the artifacts to be signed.
@@
S2, S7, S17, and S18 highlight issues surrounding the compatibility of newer signing systems with existing software systems in the software engineering process. As S18 put it, “There are teams who’ve been doing this for quite a while…Selling that is a big challenge, right? Because we need to prove that it is much more secure than what has been there in place.”
@@
S8, whose organization’s software product is built for different platforms and hardware architectures, noted compatibility issues across various architectures: “Every platform evolved separately…macOS…Windows…The requirements are changing and evolving independently”.
@@
Compatibility with other security tools can pose challenges too, as noted by S5: “The vulnerability scan report [is for]…code…CVE published…[but] signatures, I don’t think it will pick those things up.”
@@
Ease of use/Usability: Usability of the signing tool or implementation was a common challenge hampering effective use in practice. This was often expressed as developer experience, e.g., S17 said: “For a developer, make sure that they can sign the commits easily”. This point was also raised at a team level by S13: “I think that awareness of the value of signing and sophistication around how signing is performed has increased significantly, especially in recent years. But it has also resulted in a lot of smaller teams deciding that this is not a manageable thing.”
@@
With respect to documentation, S11 said: “I think there are multiple challenges there. When you build some signing infrastructure, you want to make it a self-serve so the teams can onboard themselves, how well you have your self-serve playbooks written down, documentation on using those. I think documentation can be improved.”
@@
Lack of Verification of Signatures: Lack of verification of signatures is also quoted by several subjects as a main challenge. This arises from the factors like signing serves as a “checkmark” security tool (S4: “most people who are strictly signing all commits don’t have a meaningful process for verifying those signatures or for enforcing any policy around them. They just do it because”), and lack of customer interest (S14: “We have tried to … get potential customers and customers interested…but as far as I know, none of our customers…verify our signatures on images”).
@@
Verification Of Artifact before Signing: Verification generally was highlighted as a problem, especially verifying inputs to the build process of the software delivery process. The problem noted here is that the content of software artifacts is not readily verified before signing: as S16 put it, “if you start with the component, the component gets put into a subassembly. The subassembly gets put into a package. The package can get put into a broader package or whatever. I think the people are signing. But do you know that the sum of all the pieces are accurate? … what are you really signing, and at what level?”. S11 further noted that different teams could sign arbitrary artifacts due to shared signing infrastructure.
@@ 5.2.2 Organizational Challenges
Operationalization of the Signing process: The implementations of software signing are typically neither uniform nor centralized across different organizations. Various teams are largely responsible for crafting their own solutions. As S6 said, “I can’t be more specific, but today, I’ve talked to three different teams with three different processes”. S15 suggested that automation could improve uniformity: “We try to break down the barriers with really good examples, by having common configuration of CI through similar GitHub workflows, and stuff like that. So it’s easier for people to use.”
Resource to Set up Signing: The issue of organizational resources to set up efficient signing tooling across an organization was also another cause for concern for some of our subjects. Subjects discussed this in terms of finances, time, and engineering effort. S14 said, “We need to do a lot of signing…hundreds of container images [daily]…It was a pretty significant engineering effort to build a pipeline that can handle all those signatures.”
Creating Effective Signing Policy: Policies surrounding the implementation of software signing by teams are generally not considered highly effective. These policies include elements such as role-based access, key usage, and key issuance. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S3: “… the other thing is policies, creating a policy that is, for one, it’s hard to do. And then the other part is creating an effective policy and knowing that it’s effective”
A suggested solution to address this issue is the use of regular feedback mechanisms to assess the security policies and processes of different teams within the organization. As S11 said, “We want to be more proactive…internal audits, and an assessment methodology like the NIST, I think CISF.”
No Management incentive to Sign: A lack of management incentives was also a challenge. S11 said, “The investment, the time that the engineer is putting into…signing…before a release, does it get the right attention from the senior management? Do they view it as a time well spent for the product?”
@@ 5.2.3 Human Challenges
Expertise: Some subjects noted that setting up signing tools depends on the engineer’s expertise. Unfortunately, subjects felt that most engineers lack the necessary experience and skills for this task. As S9 said, “there’s a lift there as far as…expertise”.
Developer Attitude to Signing. Software developers’ attitudes toward software signing pose challenges to the signing process. Disinterest and resistance to adopting new tools are the primary attitude-related obstacles cited. S10 observed: “Most developers don’t really understand this stuff and don’t really want to think about it…[they are] happy if they…see a big green tick[mark]”. S6 added: “They refuse to get off their old tools”
Lack of Demand from Customers: Most customer-driven organizations report that software signing is challenging to implement because customers rarely request this feature in their requirements. S14 put it like this: “Sign[ing] our images was there from day…we haven’t found many customers that are too excited about the signing aspect”.
@@ 5.3 RQ3: Perceived importance of Software Signing in mitigating risks
Finding 3. Subjects disagreed on this point. Some view signing as crucial, while others treat it as a secondary consideration in their overall strategy.
Here, we aimed to understand how important practitioners perceived software signing to be in their software delivery security. Opinions vary on the significance of signing in the software delivery process.
5.3.1 As Crucial to Provenance and Integrity
@@
Many of our subjects considered provenance an integral part of security, and felt that software signing is (or will be) a guarantee of that. S4 said, “I think signing is massively important. Signing for open source is probably the single most important thing for software supply chain security”.
5.3.2 As a Secondary Security technique

### v2

@@ 5.1.2 Build
Next, we examine our subjects’ responses regarding how signing is used in the build process of software products. For most subjects, the build process and the signing of build outputs (for subjects who do so) are automated. We expect that automation increases the use of software signing.
Verification of source code signatures before build: Signatures from source code are rarely verified before the build process begins. Four subjects mentioned some form of rigorous verification at this stage. For those who do not use or verify signatures, alternative security techniques are regularly employed, such as automating build pipelines with workflow definitions, using build policies, implementing reproducible builds, and utilizing code scanning tools to ensure the security of inputs to the build process.
Signing of build output: 14 subjects (11 organizations) reported signing at this stage of the software delivery process. 10 subjects (6 organizations) reported signing alongside attestations, SBOMS, and other metadata.
@@ 5.1.3 Package/Artifact
Verification of build signature before deployment of final artifact: Similar to the other verification steps, signatures from the build phase are rarely verified before the deployment. Only two subjects (1 organization) mentioned verification at this point. For example, S7 (Security tools, Cloud security) states, “We sign our code and then we also verify that it was built in our specific build system.”
Signing of Final Software product: All 18 subjects said that signing is done at the deployment stage to establish provenance for the organization on each of their final products.
@@ 5.1.4 Third-Party Dependencies
For third-party dependencies in the software supply chain, we assessed how software signing influences dependency selection as a security factor and how signatures are used to verify the authenticity of dependencies in practice.
@@
Presence of a Signature as a Factor for Dependency Selection: When they select third-party dependencies, most subjects do not consider whether it is signed. Only three subjects view signatures as an indicator that the package maintainer prioritizes security. Other security factors influencing the choice of third-party dependencies include the use of OpenSSF’s scorecard tool, known vulnerabilities, level of activity, licensing, and the age of the dependency.
@@
Verification of Signatures from external dependencies: Signatures on third-party dependencies are rarely verified in practice. Only three subjects (3 organizations) reported some form of verifying signatures from third-party dependencies. Signatures are mainly verified for commercial off-the-shelf software. As S11 (Firmware & testing software, Social technology) says, “If it is open source, then probably no…But if it is commercial off the shelf, we want to make sure those are signed.” However, most of our subjects reported several methods for assessing the authenticity and integrity of a third-party dependency. These methods include code and vulnerability scanning (e.g., Coverity), source rebuilding, internal mirroring of dependencies, and reliance on ecosystem internal checks (e.g., Go package manager’s checks). Other subjects perform internal reviews of dependencies. Some have a centralized dependency selection and management process, where the organization hosts and maintains dependencies (5 subjects from 4 organizations). A few subjects noted that their organization discourages or forbids third-party dependencies (S13 & S8). In addition to relying on alternative methods to assess dependency security, most of our subjects echoed the sentiment that OSS dependencies are rarely signed and that the identity of individuals signing open-source dependencies does not inherently establish trust.
@@
Signing after verification and certification of dependencies: A few subjects reported that third-party dependencies are thoroughly reviewed using one of the aforementioned techniques before being signed off for use by a team member. Six of our subjects indicated that this process is followed in their software production workflow. For example, S17 (SAAS, SSC Security tool) described their team’s process as this: “We have a signed report that’s signed by us, and we trust us…that says, ‘Hey, I scanned this third-party dependency, did all these things. I don’t have a signature for that, but I ran all these extra checks and then signed those checks. So now we have a record.”’ In contrast to that attestation approach, S18 (Security tooling, Internet service) says their team ingests third-party dependencies and signs the internal versions directly.
5.1.5 Refined Software Supply Chain Factory Model
As a result of our framework analysis, we identified various points where software signing is utilized. We compared these to the descriptions in several Software Supply Chain security frameworks, especially the SSCFM (section 4.2Figure 1).
We observed that some of these frameworks and standards are ambiguous, particularly in specifying where within the software supply chain factory model practitioners are required to establish provenance, whether by creating or verifying a signature. We therefore present a refined version of Figure 1 in Figure 3. This updated model clearly identifies the points where provenance should (could) be established, providing guidance for organizations and teams newly adopting software signing in their software creation process. For example, our refined model is compared with the SLSA framework  [79]. Despite SLSA’s rich description of security recommendations and its focus on build levels, there is surface-level confusion about whether provenance requirements apply solely to the build stage or also to earlier stages such as collaboration.
Figure 3: Our refined software supply chain factory model (compare to Figure 1), highlighting different points where provenance and integrity of software could be established (PI, PE, PB, PD, PS, PA) and Verified (VI, VE, VB, VD, VS, VA) using software signatures. These points were identified from participants’ responses. Note that while signing at PB and PA stages was commonly practiced (in green), verification at VD, VB and VE stages was minimally performed (in red). Abbreviations are defined in Table 3.
@@ 5.2 RQ2: Challenges Affecting the Implementation of Software Signing in Practice
Our next research objective was to understand the various challenges practitioners face while implementing software signing in practice. We grouped subjects’ responses under three main themes for challenges faced by practitioners in practice; Technical, Organizational, and Human factors. Table 4 gives a summary. Due to space limitations, we focus on only a subset of the individual challenges.
Table 4: Challenges to software signing implementation in practice. We have categorized related challenges highlighted by subjects. Additional technical challenges, ’Managing Signed Artifacts’ and ’Identity Management/Authentication,’ are detailed in Appendix D, as they are not exclusive to signing tools.
Observed Challenges #Subjects #Orgs Subjects’ Proposed Solutions
@@
Key Management 10 9 Use of Keyless Signing (e.g., Sigstore)
Compatibility Issues 6 6
Lack of Verification of Signatures 6 5 Signed Metadata, Component Data Management
Ease of Use/Usability 4 4 Usable Signing Tools (e.g., Sigstore), Documentation
No Unifying Standard 2 2
@@
Operationalization of the Signing Process 4 4 Automating Signing
Resources to Set up Signing 3 3
Creating Effective Signing Policy 2 2 Regular Process Feedback Mechanisms
No Management Incentive to Sign 2 2
Bureaucracy 1 1
@@
Expertise in setting setup and use of signing 5 4
Developer Attitude to Signing 3 3 Automating Signing
Lack of Demand from Customers 1 1
@@ 5.2.1 Technical Challenges
Some subjects discussed recent signing technologies that implement keyless signing (e.g., Sigstore) as a possible solution. S18 (Security tooling, Internet service) said: “Sigstore comes in…keyless signing…you don’t have to worry about long-lasting static SSH keys, or keys being compromised, or rotating keys…you don’t have these big servers to store all these signatures”. S17 (SAAS, SSC Security tool) also noted, “[Sigstore’s] other strength…I can associate my identity with an OIDC identity as opposed to necessarily needing to generate a key and keep track of that key and yada, yada. So that’s super useful because I could say, ‘Oh, this is signed with [a] GitHub identity.’ So unless my GitHub identity has been compromised, that’s much better”.
@@
Compatibility Issues: Compatibility issues arise during the integration of software signing systems and tools with other tools within the software engineering process, including other security tools and the artifacts to be signed, they also include scenarios where challenges with signing were linked to other security toolings,.
@@
S2, S7, S17, and S18 highlight issues in the compatibility of newer signing systems with existing software systems in the software engineering process. As S18 (Security tooling, Internet service) put it, “There are teams who’ve been doing this for quite a while…Selling that is a big challenge, right? Because we need to prove that it is much more secure than what has been there in place.”
@@
S8 (Internal security tool & cloud APIs, Internet services), whose organization’s software product is built for different platforms and hardware architectures, noted compatibility issues across various architectures: “Every platform evolved separately…macOS…Windows…The requirements are changing and evolving independently”.
@@
Compatibility issues can also occur in situations where current security test tools are not designed to identify issues with Software signing, as noted by S5 (Internal security tool, Cloud security): “The vulnerability scan report [is for]…code…CVE published…[but] signatures, I don’t think it will pick those things up.”
@@
Ease of Use/Usability: Usability of the signing tool or implementation was a common challenge hampering effective use in practice. This was often expressed as developer experience, e.g., S17 (SAAS, SSC Security tool) said: “For a developer, make sure that they can sign the commits easily”. This point was also raised at a team level by S13 (Internal security tool, Telecommunication): “I think that awareness of the value of signing and sophistication around how signing is performed has increased significantly, especially in recent years. But it has also resulted in a lot of smaller teams deciding that this is not a manageable thing.”
@@
Concerning documentation, S11 (Firmware & testing software, Social technology) said: “I think there are multiple challenges there. When you build some signing infrastructure, you want to make it a self-serve so the teams can onboard themselves, how well you have your self-serve playbooks written down, documentation on using those. I think documentation can be improved.”
@@
Lack of Verification of Signatures: Lack of verification of signatures is also quoted by several subjects as a main challenge. This arises from the factors like signing serves as a “checkmark” security tool (S4 (open source tooling, cloud dev tool): “most people who are strictly signing all commits don’t have a meaningful process for verifying those signatures or for enforcing any policy around them. They just do it because”), and lack of customer interest (S14 (POC Software & Security tool, cloud security): “We have tried to … get potential customers and customers interested…but as far as I know, none of our customers…verify our signatures on images.”)
@@
Verification of Artifact Before Signing: Verification generally was highlighted as a problem, especially verifying inputs to the build process of the software delivery process. The problem noted here is that the content of software artifacts is not readily verified before signing: as S16 (security tooling, Digital technology) put it, “if you start with the component, the component gets put into a subassembly. The subassembly gets put into a package. The package can get put into a broader package or whatever. I think the people are signing. But do you know that the sum of all the pieces are accurate? … what are you really signing, and at what level?”
@@ 5.2.2 Organizational Challenges
Operationalization of the Signing process: The implementations of software signing are typically neither uniform nor centralized across different organizations. Various teams are largely responsible for crafting their own solutions. As S6 (security tooling, Digital technology) said, “I can’t be more specific, but today, I’ve talked to three different teams with three different processes”. S15 (internal security tooling, Aerospace Security) suggested that automation could improve uniformity: “We try to break down the barriers with really good examples, by having common configuration of CI through similar GitHub workflows, and stuff like that. So it’s easier for people to use.”
Resource to Set up Signing: The issue of organizational resources to set up efficient signing tooling across an organization was also another cause for concern for some of our subjects. Subjects discussed this in terms of finances, time, and engineering effort. S14 (POC Software & Security tool, cloud security) said, “We need to do a lot of signing…hundreds of container images [daily]…It was a pretty significant engineering effort to build a pipeline that can handle all those signatures.”
Creating Effective Signing Policy: Policies surrounding the implementation of software signing by teams are generally not considered highly effective. These policies include elements such as role-based access, key usage, and key issuance. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S3 (SAAS, SSC Security tool): “… the other thing is policies, creating a policy that is, for one, it’s hard to do. And then the other part is creating an effective policy and knowing that it’s effective.”
A suggested solution to address this issue is the use of regular feedback mechanisms to assess the security policies and processes of different teams within the organization. As S11 (Firmware & testing software, Social technology) said, “We want to be more proactive…internal audits, and an assessment methodology like the NIST, I think CISF (NIST’s CSF).”
No Management Incentive to Sign: A lack of management incentives was also a challenge. S11 (Firmware & testing software, Social technology) said, “The investment, the time that the engineer is putting into…signing…before a release, does it get the right attention from the senior management? Do they view it as a time well spent for the product?”
@@ 5.2.3 Human Challenges
Expertise: Some subjects noted that setting up signing tools depends on the engineer’s expertise. Subjects felt that most engineers lack the necessary expertise. As S9 (SAAS, SSC security) said, “there’s a lift there as far as…expertise”.
Developer Attitude to Signing: Software developers’ attitudes toward software signing pose challenges to the signing process. Disinterest and resistance to adopting new tools are the primary attitude-related obstacles cited. S10 (SAAS, Dev tools) observed: “Most developers don’t really understand this stuff [or] want to think about it…[they are] happy if they…see a big green tick[mark].” S6 (security tooling, Digital technology) added: “They refuse to get off their old tools.”
Lack of Demand from Customers: Most customer-driven organizations report that software signing is challenging to implement because customers rarely request this feature in their requirements. S14 (POC Software & Security tool, cloud security) put it like this: “Sign[ing] our images was there from day…we haven’t found many customers that are too excited about the signing aspect.”
@@ 5.3 RQ3: Perceived Importance of Software Signing in Mitigating Risks
In this section, we aimed to understand how practitioners perceived the importance of software signing in their software delivery security. Opinions on the significance of signing in the software delivery process varied. Subjects identified three distinct perspectives on the importance of software signing, reflecting on their personal views and how signing was treated in their current and previous roles.
5.3.1 Signing as Crucial to Provenance and Integrity
@@
Many of our subjects considered provenance an integral part of security, and felt that software signing is (or will be) a guarantee of that. 10 subjects (representing 9 organizations) had the opinion that Software signing on its own is strong enough to guarantee provenance. S4 said, “I think signing is massively important. Signing for open source is probably the single most important thing for software supply chain security”. S16’s response mirrors this, “I think [Signing’s importance] is very high…There’s a senior executive VP who has to attest that the software we release is what we say it is … he could be legally sued if the software we deliver… I mean personally, not just our organization. Him, because of his level in the organization, if we publish software that we have incorrectly attested to, he can be … So it’s a very high priority.”
5.3.2 Signing as a Secondary Security Technique